What wold be the company Weighted Average Cost of Capital based on the following

Optimal proportions are 40% Debt, 10% Preferred, and 50% Common Equity Retained Earnings = $400,000 Tax = 40% Value of Cost of Debt (kd) = 10% Value of Cost of Preferred stock (Kps) = 9% Value of Cost of Equity (ke) = 14%
Answer #1

After-tax cost of debt=10*(1-tax rate)

=10*(1-0.4)=6%

WACC=Respective costs*Respective weight

=(0.4*6)+(0.1*9)+(0.5*14)

=10.3%